daddy is x years old, mum is 5 years younger than daddy and their kid is 1/2 years younger than mum. they are altogather 100 years. how old is daddy

Answer by Cromlix(4381)

Hi there,
Daddy = x
Mummy = x - 5
Kid = (x - 5)/2
x + (x - 5) + (x - 5)/2 = 100
Multiply throughout by 2
2x + 2(x - 5) + (x - 5) = 200
2x + 2x - 10 + x - 5 = 200
collect like terms:
2x + 2x + x = 200 + 10 + 5
5x = 215
x = 43
Daddy = 43
Mummy = 38
Kid = 19
Hope this helps :-)